Eric’s average income for the first 4 months of the year is $1,450.25, what must be his
average income for the remaining 8 months so that his average income for the year is
$1,780.75?
1
Expert's answer
2020-06-08T19:28:55-0400

Let, "\overline{x}_1" = Eric's average income for the first 4 months of the year = $1,450.25

"\overline{x}_2" = Eric's average income for the remaining 8 months of the year

"\overline{x}" = Eric's average income for the entire year = $1,780.75

n1 = 4 months

n2 = 8 months


Then by using the combined mean formula we have,


"\overline{x}=\frac{n_1\overline{x}_1+n_2\overline{x}_2}{n_1+n_2}"


i.e. 1780.75 = "\frac{4\times1450.25+8\overline{x}_2}{4+8}"


i.e. 5801 + 8"\overline{x}_2" = 12 x 1780.75


i.e. 8"\overline{x}_2" = 21369 - 5801 = 15568


i.e. "\overline{x}_2" = "\frac{15568}{8}" = 1946


Answer: Eric's average income for the remaining 8 months must be $1,946 so that his average income for the year will be $1,780.75.